OpenStudy (dmcdowell):

In right ∆ABC, C is a right angle and sin A = . What is mB? 52.09° 51.16° 44.32° 38.84° 37.91°

OpenStudy (dmcdowell):

@dakid88 @thuglife187 @thatone_girl @ParthKohli

OpenStudy (dakid88):

cos B = 234 / 312 B = cos^-1 ( 234 / 312) = 38.84

OpenStudy (dmcdowell):

thanks man @dakid88

OpenStudy (dakid88):

np